OpenStudy (anonymous):

the length of a rugby league field is 52 meters longer than its width "w" . The area of a rugb league field is 8160 square meters. Find the dimensions of the field

OpenStudy (anonymous):

A=8160=w*(w+52)=w^2 +52w it is a quadratic equation - solve it for w (need positive value only) let me know if you have a problem... :)

OpenStudy (anonymous):

i dont know how to solve quadratic equations

OpenStudy (anonymous):

you equation looks like: w^2 +52w-8160=0 w=[-52+/- sq.rt( 52^2-4*(8160))]/2 w=[-52+/-sq.rt(10,144)]/2 w=(-52 +/-188)/2 you choose only positive value (width can not be negative), so w=(188-52)/2 w=68 if w=68, length =68+52=120

OpenStudy (anonymous):

don't forget to put meters in your answer :)

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

welcome!